OpenStudy (anonymous):

Sawyer spent $45.22 on short-sleeve and long-sleeve T-shirts. The long-sleeve T-shirts cost $13.99 each. If he bought two of each type of shirt, what was the price of each short-sleeve T-shirt?

OpenStudy (anonymous):

Price of 2 LS shirt = 2*13.99 = 27.98 Remaining amount = 45.22-27.98=17.24 Price of SS shirt = 17.24/2 = 8.62 $

OpenStudy (anonymous):

Thank you :)